OpenStudy (camerondoherty):

Can someone help me through or show me how to do the rest of this assignment?

OpenStudy (camerondoherty):

I need to complete the chart and find the residual... but i dont even know how to find the residual...

OpenStudy (camerondoherty):

I have most of it Completed...

OpenStudy (phi):

Do you still have to fill in the chart in Part I ?

OpenStudy (phi):

I mean Question 1

OpenStudy (anonymous):

is this in flvs

OpenStudy (camerondoherty):

Yea where it asks for the y values using the equation i had

OpenStudy (phi):

you start with your equation 𝑌 = 45.547062901233 + 0.96476687061093𝑥 which I would round to y= 45.5 + 0.96x replace x with "time studying" and calculate y example: First x is 52, so figure out y = 45.55+0.965*52 = 95.73

OpenStudy (camerondoherty):

Ohhhh ok

OpenStudy (camerondoherty):

But where would i fill it in on the chart?

OpenStudy (phi):

if your calculator can save 45.547062901233 as a variable (so it's not a pain to type them all in), you can use them... otherwise, round to a reasonable number of digits)

OpenStudy (phi):

Y is the grade. its value goes in the box "grade per linear model"

OpenStudy (camerondoherty):

Ohhh I see... And when i fill it out what do do then?

OpenStudy (camerondoherty):

Do i find the differ3ence from the different y values?

OpenStudy (phi):

yes, you find the difference between the actual data and the model data (y's)

OpenStudy (phi):

the difference tells you how well the model predicts what actually happened

OpenStudy (camerondoherty):

Oh ok... Do i need the difference to calculate the residual?

OpenStudy (phi):

they tell you the sum of the differences is the residual so find all the differences ("truth" - "estimate" ) and fill in the table add them up to get the residual

OpenStudy (camerondoherty):

So there is only 1 residual?

OpenStudy (phi):

apparently.

OpenStudy (camerondoherty):

lol Thanks for all your help